OpenStudy (anonymous):

A square pyramid has a total surface area of one thousand square feet. The area of the base is two hundred square feet. What is the area of one triangular face of the lateral surface?

OpenStudy (anonymous):

thank u for looking bah

OpenStudy (bahrom7893):

What's the formula for the surface area of a square pyramid?

OpenStudy (anonymous):

\[SA=b ^{2} + 2bs\]

OpenStudy (bahrom7893):

2bs=800 bs=400

OpenStudy (bahrom7893):

what's the area of a lateral surface?

OpenStudy (anonymous):

i dont know :(

OpenStudy (anonymous):

bah???

OpenStudy (anonymous):

400?

OpenStudy (bahrom7893):

wait sorry i was away..what does a square pyramid look like again?

OpenStudy (bahrom7893):

Okay I got this

OpenStudy (bahrom7893):

bs=400 Area of a triangle is (1/2)bs=(1/2)*400=200

OpenStudy (anonymous):

the answer is 200

OpenStudy (bahrom7893):

yes